Cullumber Corporation is a regional company which is an SEC registrant. The corporation’s securities are thinly traded on NASDAQ. Cullumber Corp. has issued 20,000 units. Each unit consists of a $1,000 par, 12% subordinated debenture and 20 shares of $10 par common stock. The units were sold to outside investors for cash at $1,760 per unit. Prior to this sale, the 2-week ask price of common stock was $80 per share. Twelve percent is a reasonable market yield for the debentures, and therefore the par value of the bonds is equal to the fair value.

(a) Prepare the journal entry to record Cullumber’s transaction, under the following conditions. (Round answers to 0 decimal places, e.g. $38,487. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

(1) Employing the incremental method.
(2) Employing the proportional method, assuming the recent price quote on the common stock reflects fair value.
Answer #1

Solution

No. Account titles Debit Credit
1. Cash(1760×20000) $35,200,000
Bonds payable(1000×20000) $20,000,000
Common stock(20000*20*10) $4,000,000
Paid in capital in excess of par-Common stock $11,200,000
2. Cash $35,200,000
Bond Discount(20,000,000- 13,538,461.5) $6,461,539
Bonds payable $20,000,000
Common stock $4,000,000
Paid in capital in excess of par-Common stock(21,661,538.5**- 4000000) $17,661,539

Working note: for part 1

Cash = price received per unit * no of units = (1760×20000) = 35,200,000

bonds payable = face value * no of units = (1000×20000) = 20,000,000

common stock = no of units * 20 shares * price per share = 20000 * 20 * 10 = 4,000,000

Working note: for part 2

Allocated value:

Subordinate debentures = 1000 (as stated in question)

Common stock = no of shares * price per share = 20 * 80 = 1600

Total = common stock + debentures =1000 + 1600 = 2600

Thus ratio = 1000 / 2600 and 1600 / 2600

Amount allocate to bond = value total * ratio of bond = [35,200,000 * 10/26] = 13,538,461.5

allocated to common stock = value total * ratio of stock [35,200,000 *16/26] = 21,661,538.5

Now,

paid in capital excess of common stock = allocated common stock - common stock = 21,661,538.5 - 4,000,000 = 17,661,539

Bond discount = bonds payable - allocated to bond value = (20,000,000- 13,538,461.5) = 6,461,539